one componet of a metal sculpture consists of a solid cube with an edge of length 0.205M. The alloy used to make the cube has a density of 6.55x10^3 kg/m^3. find the cubes mass.
1
Expert's answer
2012-12-21T05:15:43-0500
M = density * V
M - mass of the cub
V - volume of the cub
density = 6.55x10^3kg/m^3
V=a^3
a - length of edge
M = density * a^3 = 6.55x10^3 kg/m^3 * (0.205m)^3 = 56kg